OpenStudy (anonymous):

what is the solution for integration sign (sinx.sec2x) dx

OpenStudy (anonymous):

\[\int\limits \frac {sinx}{\cos 2x}dx= \int\limits \frac{sinx}{2\cos^2 x-1}dx\]put cosx=y -sinxdx=dy \[\int\limits\frac{-dy}{2y^2-1}\]

OpenStudy (anonymous):

iam a big fan of urs fiz u r a great mathematician may God bless u!!!!:D